OpenStudy (anonymous):

PLEASE HELP ME WILL GIVE MEDAL! 12. If p varies directly with q, and p=10 when q=5, what is the value p when q=20? A) p=1 B) p=28 C) p=40 D) p=100 13. Simplify. Square root of 28/25 A)4square root 7 those over 5 B) square root 3 C)4square root7 those over 25 D) 2 square root 7those over 25 14. Which is the graph of the function g(x)=-|x+2|? I can't draw the graphs because I am using an iPad

OpenStudy (anonymous):

@Nurali

OpenStudy (scorcher219396):

Varies directly means y=pq So plug in your initial information y=(10)(5), y=50 Then set it up for the second situation, using what you solved y for 50=20p, p=5/2

OpenStudy (scorcher219396):

And for the second part, is the question \[\sqrt{\frac{ 28 }{ 25 }}\] or \[\frac{ \sqrt{28} }{ 25 }\] ? It makes a big difference, i'm assuming it's the first one

OpenStudy (princeharryyy):

let p = kq say k is some constant, than, 10=k.5 ................. equation1 p=k.20 .................. equation 2 divide equation1 by equation 2 to get p=40

OpenStudy (scorcher219396):

Actually, judging by the answers, it's the 25 is not under the sqrt

OpenStudy (anonymous):

It's the first one

OpenStudy (scorcher219396):

\[\frac{ \sqrt{28} }{ 25 }=\frac{ \sqrt{(4)(7)} }{ 25 }=\frac{ \sqrt{(2^2)(7)} }{ 25 }=\frac{ 2\sqrt{7} }{ 25 }\]

OpenStudy (anonymous):

Thank you @princeharryyy and @scorcher219396

OpenStudy (princeharryyy):

sqrt(28/25) = 2*sqrt(7)/5

OpenStudy (anonymous):

@scorcher219396 that is not an option though

OpenStudy (princeharryyy):

sqrt(28)/25 = 2sqrt(7)/25

OpenStudy (scorcher219396):

Yes it is, it's the last one, based on what you typed

OpenStudy (princeharryyy):

depending on what the question want to say

OpenStudy (scorcher219396):

And for the graphs, it would be a v shape (shape of absolute value), flipped upside down because of the - in front, and moved right 2 because of the -2

OpenStudy (anonymous):

Oh I'm sorry the last one is actually 2sqrt(7)/25

OpenStudy (anonymous):

No I mean it's over 5!

OpenStudy (anonymous):

Omg lol

OpenStudy (scorcher219396):

Oh gotcha. Then it would just be sqrt(25) simplified to 5

OpenStudy (scorcher219396):

(in the denominator)

OpenStudy (anonymous):

Oh okay thank you!

OpenStudy (anonymous):

I am going to opena new question so I can give u a medal lol since I already gave one to prince @scorcher219396

OpenStudy (nurali):

12. p = kq k = p/q = 10/5 = 2 p = 2q p = 2 x 20 = 40

OpenStudy (nurali):

13. Square root of 28 / 5, the square root of 28 can be reduced to the square root of 4 x 7, which reduces to 2 x square root of seven; so you end up with 2 x square root of seven / 5

OpenStudy (anonymous):

@princeharryyy I don't have that as an option

OpenStudy (anonymous):

Thank you @Nurali

OpenStudy (nurali):

Anytime.

OpenStudy (princeharryyy):

then wait a sec le' me check

OpenStudy (anonymous):

It's okay I got it hahah thank you so much though!!! <3

OpenStudy (princeharryyy):



OpenStudy (princeharryyy):

do u have this ooption

OpenStudy (princeharryyy):

option*

OpenStudy (anonymous):

Yes thank you :)

OpenStudy (princeharryyy):

welcome :)